# Draft: August 22, 2025

### v1

@@ Abstract
While governments and international organizations have set the net-zero target to prevent a climate event horizon, practical solutions are lacking mainly because of the impracticability to completely replace combustion processes. Hence, in this paper, we first design a compartmental network whose states must remain in the nonnegative orthant for physical consistency and in which the carbon dioxide emissions result from the combustion of diesel in vehicles and gas in house heaters. Then, we designed both full-state and output-feedback linear-quadratic regulators of the compartmental network to bring the mass of carbon dioxide to the pre-industrial era, which is reached in approximately 25 and 60 days, respectively. The output feedback tolerates for 6 days the combustion taking place in 5,000 vehicles and in 10,000 house heating systems, it meets the net-zero target, and it nullifies the extraction of finite natural resources. The tropospheric temperature with closed-loop reaches the equilibrium at 133 °C after 16.4 years; while such an high value requires to further investigate with climate experts the model of the dynamics of the temperature, this work is a first step in designing optimal network control systems for climate stability. Source code is publicly available [^1] https://github.com/ciroresearch/CarboNet.
Index Terms:
Circular control, circular economy, climate control.
@@ I Introduction
While governments and international organizations are adopting measures to prevent a climate event horizon, the temperature anomaly with respect to the pre-industrial era keeps increasing as recorded by NASA [1]. The main cause of this phenomenon are the emissions of greenhouse gases, especially carbon dioxide [2]. Carbon dioxide emissions result from combustion processes that, despite the enormous international effort to adopt combustion-free energy and transportation systems such as wind turbines and electric vehicles, they remain the backbone of our economy. Indeed, the share of primary energy consumption from fossil fuels in 2024 was approximately 73% in Europe and UK and 80% in China and United States [3], while the share of electric cars in use in 2024 was 4.7% in Europe, 6.4% in UK, 11% in China, and 2.7% in United States [4].
To date, the leading control community has not yet contributed significantly to Earth’s climate control as visible by searching for IEEE journal publications: as of 30 July 2025, the first author found only the works of Cavraro [5], Khargonekar et al. [6], and De Nardi et al. [7]. A recent tutorial on control theory for climate science was recently carried out by Elsherif and Taha [8], which highlights the lack of contributions from the control engineering community to Earth’s climate stabilization. Recently, Zocco et al. [9] began tackling carbon regulation as a network problem combining nonlinear control and reinforcement learning.
@@
  • Designed a compartmental network modeling the carbon emissions and the fuel consumption of a tunable number of vehicles and house heaters.
  • Designed full-state and output feedback linear-quadratic regulators of the network to bring the carbon dioxide in a target urban area to the desired pre-industrial level. While carbon control has been already addressed in Zocco et al. [9] as a network control problem, this is the first paper where the control is centralized instead of distributed at different nodes.
  • Investigated the nonlinear dynamics of the tropospheric temperature in the target urban area when the network is in closed-loop.
  • Measured the circularity of the network to both assess the consumption of natural resources and to insert the net-zero problem within the broader context of a circular economy.
The following notation is adopted throughout the paper: bold lower-case and capital letters are used for vectors and matrices, respectively; the set of nonnegative real numbers is indicated as $\overline{\mathbb{R}}_{+}$, a vector $\bm{x}\in\mathbb{R}^{n}$ with all nonnegative components is indicated as $\bm{x}\geq\geq 0$ or, equivalently, as $\bm{x}\in\overline{\mathbb{R}}^{n}_{+}$; similarly, a time-variant vector $\bm{x}(t)$ with all nonnegative components for all time is indicated as $\bm{x}(t)\in\overline{\mathbb{R}}^{n}_{+}$.
The remaining of the paper is organized as follows. Section II discusses the related work, then Section III covers the framework of linear-quadratic regulator designs along with the definition of the key variables we aim to regulate; subsequently, numerical examples and the case of $\text{CO}_{2}$ control are detailed in Section IV with full-state feedback and in Section V with output feedback; finally, Section VI concludes.
@@ II-A Circular Economy and Net Zero
The net-zero target required by the United Nations [10] and the European Commission [11] requires to drastically reduce the carbon dioxide emissions. Net-zero, essentially, means that the amount of $\text{CO}_{2}$ entering the atmosphere is equal to the amount that is removed from it [12].
The circular economy paradigm is gaining attention as a solution to decouple economic growth from sustainable use of natural resources [13]. The practices of a circular economy are, for example, reduce the material demand, reuse material for as long as possible, and repair products [14, 15]. These practices can reduce the carbon dioxide emissions because replacing the raw material extraction, transportation, and manufacturing with the less energy-intensive repair and reuse of functioning product parts and materials [16, 17, 18]. Hence, increasing circularity helps to meet the net-zero target.
The circulation of tropospheric carbon dioxide can further help to meet the net zero target mentioned earlier since carbon circulation yields a removal of carbon from the troposphere. Furthermore, carbon can be reused through photosynthesis if plants are used for carbon sequestration as discussed in Section II-B [19, 20, 21]. In this paper, we increase the circularity of tropospheric carbon dioxide via properly controlled carbon-capture systems.
@@ II-B Carbon Capture Systems
An intuitive approach to reduce the concentration of carbon dioxide in the troposphere is removing it. To this aim, different types of technologies are proposed in the literature. Rubin et al. [22] analyzed three major solutions, i.e., pre-combusion capture, post-combustion capture, and oxy-combustion capture. In power plants, the captured carbon dioxide can be piped off-shore or injected in underground locations with proper geological properties [23]. The storage of carbon dioxide in fractured rocks was investigated by Romano et al. [24].
@@
Nature-based solutions to carbon capture also exist and are divided into two macro-groups. The first group comprises the terrestrial carbon sinks such as forests, urban parks, green roofs, and green facades, while the second group comprises water bodies such as lakes and rivers [21]. In particular, the effectiveness of urban parks in the Italian city of Rome was deepened by Gratani et al. [25].
@@
Microalgae are attracting attention because of their high efficiency in carbon absorption compared to other plants and because of their use as biomass at the end of their life [19, 20]. However, most of algal studies are currently limited to laboratory conditions [26].
II-C Linear Quadratic Regulators
@@
The linear quadratic regulator (LQR) is a prominent approach for optimal control. It optimizes a positive quadratic cost function and it stabilizes linear time-invariant systems to the origin [27, 28]. Bemporad et al. [29] addressed the case of constrained system inputs and outputs, Shi et al. [30] developed finite-time horizon LQR with limited controller-system communication, while Maghfiroh et al. [31] improved the LQR performance using particle swarm optimization and the Kalman filter.
@@
Network-control based on LQR was proposed by Jaleel and Shamma [32], in which each agent computes their control actions in real-time, by Duan et al. [33] where multi-input linear systems only need to communicate with their neighbors instead of with the whole network, and by Gao et al. [34] by leveraging the theory of graphons.
@@
The major limitation of LQRs is that the choice of the weight matrices is often based on empirical considerations and it is problem dependant. To aid in this endeavor, approaches for the tuning of the weight matrices were developed by Chacko et al. [35] and by Masti et al. [36].
@@
LQRs can be designed not only in the case of access to the whole state, but also in the case of partial knowledge of the state. Solutions to the latter situation were provided by Levine and Athans [37], by Nersesov et al. [27], and by Ilka and Murgovski [28].
@@
Examples of applications of the LQR are wind turbine vibration control [38], control of a single-phase inverter [39], and optimal drug delivery [27]. In this paper, we develop both output and full-state feedback LQRs for network-based carbon control.
@@ III Network, Key Variables, and Linear-Quadratic Regulators
This section is divided into three parts: the design of the compartmental network is covered in the first part, then the key sustainability variables to be controlled are introduced, and finally, the LQR design procedure is detailed.
@@ III-A Network Design
The compartmental diagraph for carbon control considered in this paper is depicted in Fig. 1 and it is based on the formalism of thermodynamical material networks [40, 41, 42, 43, 44]. Specifically, the green node ($c^{5}_{5,5}$) is the “CC system”, which is the abbreviation for “carbon capture system”. Its role is to remove the $\text{CO}_{2}$ from the red node ($c^{1}_{1,1}$), which indicates the urban area whose carbon dioxide we aim to regulate via the control laws covered in Section III. The tropospheric $\text{CO}_{2}$ in the urban area is increased by two types of sources: by $n_{\text{q}}$ vehicles with internal-combustion engines ($c^{2}_{2,2}$) and by $n_{\text{h}}$ heaters ($c^{3}_{3,3}$); the sources of carbon take and burn the fuel from the natural reserve of oil or gas ($c^{6}_{6,6}$). The urban area exchanges carbon with the surrounding area indicated with $c^{4}_{4,4}$. Overall, the system has the four states $x_{i}(t)$ indicated in blue in Fig. 1, with $i\in\{1,2,3,4\}$, and the control signal $u(t)$ indicated in orange. The states $x_{1}(t)$ and $x_{4}(t)$ are masses of $\text{CO}_{2}$ in the troposphere, while the states $x_{2}(t)$ and $x_{3}(t)$ are masses of fuel converted into $\text{CO}_{2}$ via combustion to move the $n_{\text{q}}$ vehicles (in $c^{2}_{2,2}$) or to heat-up the $n_{\text{h}}$ buildings (in $c^{3}_{3,3}$).
@@
The system in Fig. 1 can be represented as the thermodynamical material network [40, 41, 42, 43, 44]
@@
With the network (1), it holds that the number of compartments, the number of nodes, and the number of arcs are $n_{\text{c}}=15$, $n_{\text{v}}=6$, and $n_{\text{a}}=9$, respectively.
@@
$\begin{aligned} \dot{x}_{4}(t)=a_{41}x_{1}(t)-a_{14}x_{4}(t)+n_{\text{q}}a_{42}x_{2}(t)\\ +n_{\text{h}}a_{43}x_{3}(t),\,\,x_{4}(0)=x_{40}.\end{aligned}$ (6)
From (3)-(6), we have that the state-space matrix form of the network in Fig. 1 is
@@
$\bm{A}=\begin{bmatrix}-a_{41}&n_{\text{q}}a_{12}&n_{\text{h}}a_{13}&a_{14}\\ 0&\theta_{1}&0&0\\ 0&0&\theta_{2}&0\\ a_{41}&n_{\text{q}}a_{42}&n_{\text{h}}a_{43}&-a_{14}\end{bmatrix},\bm{B}=\begin{bmatrix}-1\\ 0\\ 0\\ 0\end{bmatrix}$, (9)
where $\theta_{1}=-n_{\text{q}}(a_{12}+a_{42})+a_{22}$ and $\theta_{2}=-n_{\text{h}}(a_{13}+a_{43})+a_{33}$. The dimension $l$ is the number of measured states. In Section III-C we will cover the case of full-state feedback (i.e., $l=n=4$), while Section III-D will cover the case of output feedback (i.e., $l<n=4$).
@@ III-B1 Circularity of $\text{CO}_{2}$
Thus, the instantaneous circularity is [42]
$\lambda(\mathcal{N};t)=-\left(\Delta\mu_{\text{fb}}\overline{m}_{\textup{f}\textup{b}}(t)+\dot{\overline{m}}_{\textup{f}\textup{c}}(t)\right).$ (10)
In (10), $\mathcal{N}$ is a thermodynamical material network [42, 40, 41, 44, 43], $m_{\text{f}\text{b}}(t)$ is the net finite-time sustainable mass transported in batches (e.g., solids transported on trucks) and $\dot{m}_{\text{f}\text{c}}(t)$ is the net finite-time sustainable flow transported continuously (e.g., fluids transported through pipes). The quantity
$\overline{m}_{\text{fb}}(t)=c_{\text{fb}}m_{\text{fb}}(t)$ (11)
is the weighted $m_{\text{fb}}(t)$, where $c_{\text{fb}}\in(0,1]$ is the criticality coefficient of the material indicating to what extent the material is critical (e.g., high value for a critical raw material [46]), and where $\mu_{\text{fb}}\geq 1$ takes into account how many batches of finite-time sustainable materials are disposed of while they are still functional (i.e., disposing of faulty parts is better than disposing of working parts). Similarly, $\dot{\overline{m}}_{\text{fc}}(t)=c_{\text{fc}}\dot{m}_{\text{fc}}(t)$ is the weighted $\dot{m}_{\text{fc}}(t)$, where $c_{\text{fc}}\in[1,+\infty)$ is the criticality coefficient of the finite-time sustainable material transported continuously. Finally, $\Delta>0$ is a constant with the unit of a frequency introduced as a multiplying factor in order to convert the mass $\overline{m}_{\text{f}\text{b}}(t)$ into a flow, and hence, to make the sum with $\dot{\overline{m}}_{\text{f}\text{c}}(t)$ physically consistent. The choice of $\Delta$ is arbitrary, but its value must be kept the same for any calculation of $\lambda(\mathcal{N};t)$ in order to make comparisons. Thus, $\lambda(\mathcal{N};t)\in(-\infty,+\infty)$ and the circularity problem can be stated as
@@
For the case of $\mathcal{N}_{\text{c}}$ in Fig. 1, it holds that $m_{\text{fb}}(t)=0$. With $c_{\text{fc}}$ = 1 due to the high criticality of tropospheric $\text{CO}_{2}$, the instantaneous circularity (10) yields
@@
Note that, in line with Definition 1, $\phi_{1}(t)$ takes into account the material flows extracted from non-renewable reserves (i.e., $c^{6}_{6,6}$), while $\phi_{\text{nz}}(t)$ is the difference between the mass flow rate of $\text{CO}_{2}$ emitted to the atmosphere (which is considered a pollutant because well beyond the natural atmospheric concentrations) and the mass flow rate of $\text{CO}_{2}$ removed by the carbon-capture system, i.e., $u(t)$. Specifically, $\phi_{\text{nz}}(t)$ corresponds to the net-zero target and it is met when
@@
where $\gamma\in\overline{\mathbb{R}}_{+}$ is a constant. The constant $\gamma$ is the concentration of $\text{CO}_{2}$ in the considered part of the troposphere, i.e., in the compartments $c^{1}_{1,1}$ and $c^{4}_{4,4}$ together, that in the net-zero condition have the total input flow equal to the total output flow (i.e., $u(t)$). Note that the exchange of carbon inside the troposphere, i.e., between $c^{1}_{1,1}$ and $c^{4}_{4,4}$, is not constrained in the net-zero target, and hence, in general, $x_{1}(t)$ and $x_{4}(t)$ can vary over time even if condition (16) is reached.

### v2

# CarboNet: A Finite-Time Combustion-Tolerant Compartmental Network for Tropospheric Carbon Control

@@ Abstract
While governments and international organizations have set the net-zero target to prevent a climate event horizon, practical solutions are lacking mainly because of the impracticability in completely replacing combustion processes. To address the net-zero target problem, in this paper we first design a compartmental network whose states must remain in the nonnegative orthant for physical consistency and in which the carbon dioxide emissions result from the combustion of diesel in vehicles and gas in house heaters. Then, we design both full-state and output-feedback linear-quadratic regulators of the compartmental network to bring the mass of carbon dioxide to the pre-industrial era, which is reached in approximately 25 and 60 days, respectively. The output-feedback controller tolerates for 6 days the combustion taking place in 5,000 vehicles and in 10,000 house heating systems, meets the net-zero target, and nullifies the extraction of finite natural resources. With closed-loop control, the tropospheric temperature stabilizes approximately to the pre-industrial era reference condition, i.e., to 13.5 °C, which is 21.7 °C lower than the steady-state temperature achieved without carbon capture. This work is a first step in designing optimal network control systems for climate stability. Source code is publicly available [^1] https://github.com/ciroresearch/CarboNet.
@@ I Introduction
While governments and international organizations are adopting measures to prevent a climate event horizon, the temperature anomaly with respect to the pre-industrial era keeps increasing as recorded by NASA [NASAtemperature]. The main cause of this phenomenon are the emissions of greenhouse gases, especially carbon dioxide [NASAcarbon]. Carbon dioxide emissions result from combustion processes that are still the backbone of our economy despite the enormous international effort to adopt combustion-free energy and transportation systems, such as wind turbines and electric vehicles. Specifically, the share of primary energy consumption from fossil fuels in 2024 was approximately 73% in Europe and the UK and 80% in China and the United States [DatafossilEnergy], while the share of electric cars in use in 2024 was 4.7% in Europe, 6.4% in the UK, 11% in China, and 2.7% in the United States [DataElecVehicles].
To the best of our knowledge, the analysis and control of Earth’s climate via the methodological tools of dynamical systems and control theory is not widely explored as visible by searching for IEEE journal publications: as of 30 July 2025, the authors found only the works of Cavraro [cavraro2024advocating], Khargonekar et al. [khargonekar2024climate], and De Nardi et al. [de2025climate]. A recent tutorial on control theory for climate science was recently carried out by Elsherif and Taha [elsherif2025climate], which highlights the lack of contributions from the control engineering community to the Earth’s climate stabilization. Recently, Zocco et al. [zocco2025circularCarbon] began addressing carbon regulation as a network problem combining nonlinear control and reinforcement learning.
@@
  • Developes a compartmental network model for carbon emissions and fuel consumption of a tunable number of vehicles and house heaters.
  • Designes full-state and output feedback linear-quadratic regulators (LQRs) for the network to bring the carbon dioxide level in a target urban area to the pre-industrial reference value. While carbon control has been already addressed in Zocco et al. [zocco2025circularCarbon] as a network control problem, this is the first paper where the control is centralized instead of distributed at different nodes.
  • Investigates the nonlinear dynamics of the tropospheric temperature in the target urban area for the closed-loop network.
  • Measures the circularity of the network to assess the consumption of natural resources and to insert the net-zero problem within the broader context of a circular economy.
The following notation is adopted throughout the paper. Bold lower-case and capital letters are used for vectors and matrices, respectively; the set of nonnegative real numbers is denoted by $\overline{\mathbb{R}}_{+}$, a vector $\bm{x}\in\mathbb{R}^{n}$ with all nonnegative components is denoted by $\bm{x}\geq\geq 0$ or, equivalently, by $\bm{x}\in\overline{\mathbb{R}}^{n}_{+}$; similarly, a time-variant vector $\bm{x}(t)$ with all nonnegative components for all time is denoted by $\bm{x}(t)\in\overline{\mathbb{R}}^{n}_{+}$, $t\geq 0$.
The remainder of the paper is organized as follows. Section II discusses related work, while Section III covers the framework of linear-quadratic regulator designs along with the definition of the key variables we aim to regulate. Numerical examples and the case of $\text{CO}_{2}$ control are detailed in Section IV for full-state feedback control and Section V for output feedback control. Finally, in Section VI we provide conclusions.
@@ II-A Circular Economy and Net Zero
The net-zero target required by the United Nations [UN-netZero] and the European Commission [EU-netZero] requires to drastically reduce the carbon dioxide emissions. Net-zero essentially means that the amount of $\text{CO}_{2}$ entering the atmosphere is equal to the amount that is removed from it [NZdefinition].
The circular economy paradigm is gaining attention as a solution to decouple economic growth from sustainable use of natural resources [alexander2023handbook]. The practices of a circular economy include reducing the material demand, reusing material for as long as possible, and repairing products [potting2017circular, velenturf2021principles]. These practices can reduce the carbon dioxide emissions by replacing the raw material extraction, transportation, and manufacturing with the less energy-intensive repair and reuse of functioning product parts and materials [bonsu2020towards, yu2021evaluating, marche2022comparison]. Hence, increasing circularity helps meet the net-zero target.
The circulation of tropospheric carbon dioxide can further help meet the net-zero target mentioned above since carbon circulation yields a removal of carbon from the troposphere. Furthermore, carbon can be reused through photosynthesis if plants are used for carbon sequestration as discussed in Section II-B and in [zhou2017bio, tsai2017potential, pereira2024nature]. In this paper, we increase the circularity of tropospheric carbon dioxide via properly controlled carbon-capture systems.
@@ II-B Carbon Capture Systems
An intuitive approach to reduce the concentration of carbon dioxide in the troposphere is removing it. To this aim, different types of technologies are proposed in the literature. Rubin et al. [rubin2012outlook] analyzed three major solutions, namely, pre-combusion capture, post-combustion capture, and oxy-combustion capture. In power plants, the captured carbon dioxide can be piped off-shore or injected in underground locations with proper geological properties [wilberforce2021progress]. The storage of carbon dioxide in fractured rocks was investigated by Romano et al. [romano2025evaluation].
@@
Nature-based solutions to carbon capture also exist and are divided into two macro-groups. The first group comprises of terrestrial carbon sinks such as forests, urban parks, green roofs, and green facades, while the second group comprises of water bodies such as lakes and rivers [pereira2024nature]. In particular, the effectiveness of urban parks in the Italian city of Rome was deepened by Gratani et al. [gratani2016carbon].
@@
Microalgae are attracting attention because of their high efficiency in carbon absorption compared to other plants and because of their use as biomass at the end of their life [zhou2017bio, tsai2017potential]. However, most of algal studies are currently limited to laboratory conditions [yang2024carbon]. In this paper, we design a network involving a carbon sink and find an optimal absorption rate of carbon to recover the pre-industrial era reference condition. We do not assume any particular carbon capture system to perform the optimal absorption rate.
II-C Linear-Quadratic Regulators
@@
The linear-quadratic regulator (LQR) is a prominent approach for optimal control. It optimizes a positive quadratic cost function and it stabilizes linear time-invariant as well as time-varying systems to the origin [nersesov2004optimal, ilka2022novel]. Bemporad et al. [bemporad2002explicit] addressed the case of constrained system inputs and outputs, Shi et al. [shi2012finite] developed finite-time horizon LQR with limited controller-system communication, while Maghfiroh et al. [maghfiroh2022improved] improved the LQR performance using particle swarm optimization and the Kalman filter.
@@
Network-control based on LQR was proposed by Jaleel and Shamma [jaleel2017design], in which each agent computes their control actions in real-time, by Duan et al. [duan2022distributed] where multi-input linear systems only need to communicate with their neighbors instead of with the whole network, and by Gao et al. [gao2021subspace] by leveraging the theory of graphons.
@@
The major limitation of LQRs is that the choice of the weighting matrices is often based on empirical considerations and is problem dependant. To aid in this endeavor, approaches for the tuning of the weighting matrices were developed by Chacko et al. [chacko2024optimizing] and by Masti et al. [masti2021tuning].
@@
LQRs can be designed not only in the case of access to the whole state, but also in the case of partial knowledge of the state. Solutions to the latter situation were provided by Levine and Athans [levine2003determination], by Nersesov et al. [nersesov2004optimal], and by Ilka and Murgovski [ilka2022novel].
@@
Examples of applications of the LQR are wind turbine vibration control [wang2024filter], control of a single-phase inverter [arab2019lqr], and optimal drug delivery [nersesov2004optimal]. In this paper, we develop both output and full-state feedback LQRs for network-based carbon control.
@@ III Network, Key Variables, and Linear-Quadratic Regulators
This section is divided into three parts. In the first part, the design of the compartmental network is covered, then the key sustainability variables to be controlled are introduced, and finally, the LQR design procedure is detailed.
@@ III-A Network Design
The compartmental diagraph for carbon control considered in this paper is depicted in Fig. 1 and it is based on the formalism of thermodynamical material networks [zocco2023thermodynamical, zocco2025synchronized, zocco2025cirl, zocco2022circularity, zocco2025circular]. Specifically, the green node ($c^{5}_{5,5}$) is the carbon capture (CC) system. Its role is to remove the $\text{CO}_{2}$ from the red node ($c^{1}_{1,1}$), which indicates the urban area whose carbon dioxide we aim to regulate via the control laws covered in Section III. The tropospheric $\text{CO}_{2}$ in the urban area is increased by two types of sources: by $n_{\text{q}}$ vehicles with internal-combustion engines ($c^{2}_{2,2}$) and by $n_{\text{h}}$ heaters ($c^{3}_{3,3}$); the sources of carbon take and burn the fuel from the natural reserve of oil or gas ($c^{6}_{6,6}$). The urban area exchanges carbon with the surrounding area, which is indicated by $c^{4}_{4,4}$. Overall, the system has the four states $x_{i}(t)$ indicated in blue in Fig. 1, with $i\in\{1,2,3,4\}$, and the control signal $u(t)$ is indicated in orange. The states $x_{1}(t)$ and $x_{4}(t)$ are masses of $\text{CO}_{2}$ in the troposphere, while the states $x_{2}(t)$ and $x_{3}(t)$ are masses of fuel converted into $\text{CO}_{2}$ via combustion to move the $n_{\text{q}}$ vehicles (in $c^{2}_{2,2}$) or to heat-up the $n_{\text{h}}$ buildings (in $c^{3}_{3,3}$).
@@
The system in Fig. 1 can be represented as the thermodynamical material network (see [zocco2023thermodynamical, zocco2025synchronized, zocco2025cirl, zocco2022circularity, zocco2025circular])
@@
For the network (1), the number of compartments, the number of nodes, and the number of arcs are $n_{\text{c}}=15$, $n_{\text{v}}=6$, and $n_{\text{a}}=9$, respectively.
@@
$\begin{aligned} \dot{x}_{4}(t)=a_{41}x_{1}(t)-a_{14}x_{4}(t)+n_{\text{q}}a_{42}x_{2}(t)\\ +n_{\text{h}}a_{43}x_{3}(t),\,\,x_{4}(0)=x_{40},\end{aligned}$ (6)
which in state-space form is given by
@@
$\bm{A}=\begin{bmatrix}-a_{41}&n_{\text{q}}a_{12}&n_{\text{h}}a_{13}&a_{14}\\ 0&\theta_{1}&0&0\\ 0&0&\theta_{2}&0\\ a_{41}&n_{\text{q}}a_{42}&n_{\text{h}}a_{43}&-a_{14}\end{bmatrix},\quad\bm{B}=\begin{bmatrix}-1\\ 0\\ 0\\ 0\end{bmatrix}$, (9)
where $\theta_{1}=-n_{\text{q}}(a_{12}+a_{42})+a_{22}$ and $\theta_{2}=-n_{\text{h}}(a_{13}+a_{43})+a_{33}$. The dimension $l$ is the number of measured states. In Section III-C we will cover the case of full-state feedback (i.e., $l=n=4$), while Section III-D will cover the case of output feedback (i.e., $l<n=4$). The matrix $\bm{C}$ will take different forms in those two sections.
@@ III-B1 Circularity of $\text{CO}_{2}$
Thus, the instantaneous circularity is given by [zocco2025cirl]
$\lambda(\mathcal{N};t)=-\left(\Delta\mu_{\text{fb}}\overline{m}_{\textup{f}\textup{b}}(t)+\dot{\overline{m}}_{\textup{f}\textup{c}}(t)\right),$ (10)
where $\mathcal{N}$ is a thermodynamical material network [zocco2025cirl, zocco2023thermodynamical, zocco2025synchronized, zocco2025circular, zocco2022circularity], $m_{\text{f}\text{b}}(t)$ is the net finite-time sustainable mass transported in batches (e.g., solids transported on trucks), and $\dot{m}_{\text{f}\text{c}}(t)$ is the net finite-time sustainable flow transported continuously (e.g., fluids transported through pipes). The quantity $\overline{m}_{\text{fb}}(t)=c_{\text{fb}}m_{\text{fb}}(t)$ is the weighted $m_{\text{fb}}(t)$, where $c_{\text{fb}}\in(0,1]$ is the criticality coefficient of the material indicating to what extent the material is critical (e.g., high value for a critical raw material [CRMs-EU]), and $\mu_{\text{fb}}\geq 1$ takes into account how many batches of finite-time sustainable materials are disposed of while they are still functional (i.e., disposing of faulty parts is better than disposing of working parts). Similarly, $\dot{\overline{m}}_{\text{fc}}(t)=c_{\text{fc}}\dot{m}_{\text{fc}}(t)$ is the weighted $\dot{m}_{\text{fc}}(t)$, where $c_{\text{fc}}\in(0,1]$ is the criticality coefficient of the finite-time sustainable material transported continuously. Finally, $\Delta>0$ is a constant with the unit of a frequency introduced as a multiplying factor in order to convert the mass $\overline{m}_{\text{f}\text{b}}(t)$ into a flow, and hence, makes the sum with $\dot{\overline{m}}_{\text{f}\text{c}}(t)$ physically consistent. The choice of $\Delta$ is arbitrary, but its value must be kept the same for any calculation of $\lambda(\mathcal{N};t)$ in order to make meaningful comparisons. Thus, $\lambda(\mathcal{N};t)\in(-\infty,+\infty)$ and the circularity problem can be stated as
@@
For $\mathcal{N}_{\text{c}}$ shown in Fig. 1, note that $m_{\text{fb}}(t)=0$. For $c_{\text{fc}}$ = 1, due to the high criticality of tropospheric $\text{CO}_{2}$, the instantaneous circularity (10) yields
@@
Note that, consistent with Definition 1, $\phi_{1}(t)$ takes into account the material flows extracted from non-renewable reserves (i.e., $c^{6}_{6,6}$), whereas $\phi_{\text{nz}}(t)$ is the difference between the mass flow rate of $\text{CO}_{2}$ emitted to the atmosphere (which is considered a pollutant because it is well beyond the natural atmospheric concentrations) and the mass flow rate of $\text{CO}_{2}$ removed by the carbon-capture system, i.e., $u(t)$. Specifically, $\phi_{\text{nz}}(t)$ corresponds to the net-zero target and it is met when
@@
where $\gamma\in\overline{\mathbb{R}}_{+}$ is a constant. The constant $\gamma$ is the mass of $\text{CO}_{2}$ in the considered part of the troposphere, i.e., in the compartments $c^{1}_{1,1}$ and $c^{4}_{4,4}$ together, that, for a net-zero condition, have the total input flow equal to the total output flow (i.e., $u(t)$). Note that the exchange of carbon inside the troposphere, i.e., between $c^{1}_{1,1}$ and $c^{4}_{4,4}$, is not constrained in the net-zero target, and hence, in general $x_{1}(t)$ and $x_{4}(t)$ can vary over time even if condition (15) is reached.